OpenStudy (anonymous):

which of the following is NOT equivelent to the rational expression: 4/3-x

OpenStudy (mertsj):

Where is "the following"?

OpenStudy (anonymous):

a. -4/x-3 b.-4/x-3 c.-4/3-x

OpenStudy (mertsj):

Choices a and b are exactly the same.

OpenStudy (anonymous):

a has a -4, and for b the negative is for the whole problem:)

jimthompson5910 (jim_thompson5910):

4/(3-x) = 4/(-(x-3)) = -4/(x-3) So 4/(3-x) is equivalent to -4/(x-3) which means that choices A and B are out. So choice C is the answer.